# MELA-TTS: JOINT TRANSFORMER-DIFFUSION MODEL WITH REPRESENTATION ALIGNMENT FOR SPEECH SYNTHESIS

### v1

@@ 1 Introduction
Fig. 1: The joint transformer and diffusion architecture. The autoregressive transformer decoder generates continuous vectors $\bf{h}$ as the condition to the diffusion model to generate the mel-spectrogram chunk.
@@ 2.2 Diffusion for mel-spectrogram generation
Here $\Psi_{i}=\{[h_{i-1},h_{i}],{\bf v},{\bf u}\}$ is the condition, and ${\bf X}^{(i)}_{t}=\alpha_{t}{\bf X}^{(i)}_{0}+\sigma_{t}\epsilon$ is given by a diffusion forward process [8], and $\epsilon$ is the standard gaussian noise. We follow the variance preserving (VP) formulation and set $\alpha_{t}=\rm{cos}(\frac{\pi t}{2})$ and $\sigma_{t}=\rm{sin}(\frac{\pi t}{2})$. The previous continuous vector $h_{i-1}$ and the mel-spectrogram chunk ${\bf X}^{(i-1)}_{0}$ are provided as prefix context for the diffusion model, and the output of the prefix part will be discarded. The loss is defined as the L2 distance between the predicted mel-spectrogram and the ground truth mel-spectrogram:
@@ 2.3 Representation alignment module
For alignment objective, one might intuitively consider using the mel-spectrogram directly as the alignment target. However, experimental results revealed that this strategy fails to provide any positive gains and instead significantly degrades both content consistency and speaker similarity in voice cloning. We will discuss it in section 3.2.
@@ 3.1.2 Model configuration
In MELA-TTS, all waveforms are resampled at 24 kHz, and the feature is the 80-dimensional mel-spectrogram extracted at 50 Hz with a window length of 1920 and a hop length of 480. The default size of the mel chunk is 8 (160ms). Thus, the autoregressive transformer works at a rate of 6.25 Hz (50/8 Hz) to generate continuous vectors $\bf{h}$, much smaller than most discrete-token-based TTS systems (25Hz for CosyVoice3 and 75Hz for VALL-E). With the interleaving ratio $n:m$ set to 4:3, MELA-TTS generates 3 mel-spectrogram chunks (480ms of speech) for every 4 text tokens received.
The transformer decoder follows the configuration of the pre-trained textual LLM, Qwen2-0.5B [15], and is initialized using its weights. The diffusion module is a 22-layer diffusion transformer with 1024 hidden states and 16 heads. Following [16], the diffusion module is trained on both conditional and non-conditional situations to enable the classifier-free guidance (CFG) [17] at inference:
@@ 3.1.3 Metrics
We evaluate MELA-TTS’s speech generation with CER/WER for content consistency, and cosine similarity between generated speech and reference speech for voice cloning speaker similarity (SS). Specifically, we use Whisper-large V3 [20] to calculate English WER and Paraformer [21] to calculate Chinese CER. SS is calculated on the speaker embedding extracted by WavLM-TDNN [22] (the result is denoted as SS1), or the ERes2Net speaker verification model [23] (denoted as SS2).
@@ 3.2 Ablation study on LibriTTS
Table 1 further compares the performance of streaming and offline synthesis. In streaming mode, MELA-TTS exhibits comparable WER to offline mode: 6.3 vs. 6.6 for the baseline condition, and 5.2 vs. 5.0 when both the representation alignment and utterance embedding modules are incorporated. Furthermore, both SS1 and SS2 remain nearly identical to those obtained under the offline configuration, which demonstrates great robustness of MELA-TTS in streaming mode.
@@ 3.3 Evaluation on large-scale data
For voice cloning speaker similarity, MELA-TTS with representation is comparable with competitive models on tesh-zh and test-hard, but lags in test-en. A possible reason for the suboptimal performance on speaker similarity is that in MELA-TTS, the diffusion module can only leverage the local context, while in discrete-token-based multi-stage system, like CosyVoice series, the diffusion or flow-matching module can utilize all input tokens (all history tokens in the streaming mode), as well as the prompt speech, as conditions to generate the mel-spectrogram. Similar speaker similarity gaps have also been observed in other continuous representations-based systems [3, 24]. We leave the optimization of the voice cloning ability for future work.

### v2

@@ 1 Introduction
Fig. 1: The joint transformer and diffusion architecture. The autoregressive transformer decoder generates continuous vectors $\bf{h}$ as the condition to the diffusion model to generate the mel chunks.
@@ 2.2 Diffusion for mel-spectrogram generation
Here $\Psi_{i}=\{[h_{i-1},h_{i}],{\bf v},{\bf u}\}$ is the condition, and ${\bf X}^{(i)}_{t}=\alpha_{t}{\bf X}^{(i)}_{0}+\sigma_{t}\epsilon$ is given by a diffusion forward process [18], and $\epsilon$ is the standard gaussian noise. We follow the variance preserving (VP) formulation and set $\alpha_{t}=\rm{cos}(\frac{\pi t}{2})$ and $\sigma_{t}=\rm{sin}(\frac{\pi t}{2})$. The previous continuous vector $h_{i-1}$ and the mel-spectrogram chunk ${\bf X}^{(i-1)}_{0}$ are provided as prefix context for the diffusion model, and the output of the prefix part will be discarded. The loss is defined as the L2 distance between the prediction and the ground truth mel-spectrogram:
@@ 2.3 Representation alignment module
For alignment objective, one might intuitively consider using mel-spectrogram directly as the alignment target. However, experimental results revealed that this strategy fails to provide positive gains and instead significantly degrades both content consistency and speaker similarity in voice cloning. We will discuss it in 3.2.
@@ 3.1.2 Model configuration
In MELA-TTS, waveforms are resampled at 24 kHz, and the feature is 80-dimensional mel-spectrogram extracted at 50 Hz with a window length of 1920 and a hop length of 480. The size of the mel chunk is 8 (160ms). Thus, the autoregressive transformer works at a rate of 6.25 Hz (50/8 Hz) to generate continuous vectors $\bf{h}$, much smaller than most discrete-token-based TTS systems (25Hz for CosyVoice3 and 75Hz for VALL-E). With the interleaving ratio $n:m=$ 4:3, MELA-TTS generates 3 mel-spectrogram chunks for every 4 text tokens. A HiFTNet-based vocoder [14] is used to reconstruct the waveform from the mel-spectrograms.
The transformer decoder follows the configuration of the pre-trained textual LLM, Qwen2-0.5B [20], and is initialized using its weights. The diffusion module is a 22-layer diffusion transformer with 1024-dimensional hidden states and 16 heads. Following [12], the diffusion module is trained on both conditional and non-conditional situations to enable the classifier-free guidance (CFG) [10] at inference:
@@ 3.1.3 Metrics
We evaluate MELA-TTS using CER/WER for content consistency, and cosine similarity between generated speech and reference speech for voice cloning speaker similarity (SS). Specifically, we use Whisper-large V3 [19] to calculate English WER and Paraformer [9] for Chinese CER. SS is calculated on the speaker embedding extracted by WavLM-TDNN [3] (the result is denoted as SS1), or the ERes2Net speaker verification model [4] (denoted as SS2). For subjective evaluation, we conducted an additional A/B preference test with 15 participants, each of whom was asked to complete 15 paired comparisons of speech samples.
@@ 3.2 Ablation study on LibriTTS
Table 1 compares the performance of streaming and offline synthesis. In streaming mode, MELA-TTS exhibits comparable WER to offline mode: 6.3 vs. 6.6 for the baseline condition, and 5.2 vs. 5.0 when representation alignment and utterance embedding modules are incorporated. Furthermore, both SS1 and SS2 remain nearly identical to those obtained under the offline configuration, which demonstrates great robustness of MELA-TTS in streaming mode.
@@ 3.3 Evaluation on large-scale data
For voice cloning, MELA-TTS is comparable with competitive models on tesh-zh and test-hard, but lags in test-en. A possible reason for the suboptimal performance on speaker similarity is that in MELA-TTS, the diffusion module can only leverage the local context, while in discrete-token-based multi-stage system, like CosyVoice series, the diffusion or flow-matching module can utilize all input tokens (all history tokens in the streaming mode), as well as the prompt speech, as conditions to generate the mel-spectrogram. Similar speaker similarity gaps have also been observed in other continuous-representation-based systems [15, 23]. We leave the optimization of the voice cloning ability for future work.
Fig. 5: Subjective preference between MELA-TTS and CosyVoice.
The results of subjective evaluation are shown in Figure 5. The proposed method is preferred over CosyVoice2 and CosyVoice3 in 66.7% and 57.3% of the trials, indicating better perceptual quality than discrete-token based counterparts.